# Sparse Representations of Dynamical Networks:

### v4

@@ 3.4 System Response-Type Realizations
Let the system described by the realization (4a)-(4b), of order $n$, have a TFM be denoted $\mathbf{G}(\lambda)$. Then, a pair $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$ of proper TFMs, where $\mathbf{W}(\lambda)\in\mathbb{R}^{p\times p}(\lambda)$ and $\mathbf{V}(\lambda)\in\mathbb{R}^{p\times m}(\lambda)$, with $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ having a McMillan degree of at most $n-p$ and which satisfy
@@
It is worthwhile to point out that the two TFMs which make up the SRTR pair denoted $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$ fulfill the same role as the TFMs $\widetilde{\mathbf{R}}^{+}(z)$ and $\widetilde{\mathbf{N}}(z)$ from the state iteration equation of (18) from [35]. Denoting by $\mathbf{B}(z)$ the $\mathcal{Z}$-transform of the state vector belonging to the controller, the aforementioned state iteration is expressed in the same manner as (8) via the identity
@@
It is this very fact which brought about the designation of System Response-Type Realization and which can be exploited to obtain sparse and scalable representations for systems in the context of distributed control. Note that, in contrast to the implementations from [35], the one from (8) is also available for continuous-time systems.
@@
In the sequel, SRTR pairs of controllers will be employed to implement control laws in a distributed fashion. However, in contrast to the approach from [35] and its implementations of type (14), we point out that closed-loop maps are not intrinsically linked to the definition of an SRTR pair. Therefore, the newly introduced representation bears all the hallmarks of the so-called implementation matrices from [13], in which a mutation of the classical SLS paradigm allows for the decoupling of controller synthesis from closed-loop design. Although the framework proposed in the sequel shares this benefit, we continue to refer to SRTR pairs as such due to (14) being the inspiration for the proposed representation and, more importantly, due to not wishing to overshadow their significant potential for system analysis, which would inevitably result from naming them “implementation matrices”.
@@
Another feature of the SRTR representation is that any pair $\big{(}\mathbf{W}(\lambda),$ $\mathbf{V}(\lambda)\big{)}$ possesses the same sparsity pattern with its subsequent NRF pair $\big{(}\mathbf{\Phi}(\lambda),\mathbf{\Gamma}(\lambda)\big{)}$, obtained via (9a)-(9b). Therefore, $\mathbf{W}(\lambda)$ is lower triangular if and only if $\mathbf{\Phi}(\lambda)$ is lower triangular and, similarly, $\mathbf{V}(\lambda)$ is tridiagonal if and only if $\mathbf{\Gamma}(\lambda)$ is tridiagonal.
@@ 1. $\mathbf{a)}$
(15) $\small\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\begin{array}[]{c|cc}A_{22}+KA_{12}-\lambda I_{n-p}&KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\\ \hline\cr A_{12}&A_{11}-A_{12}K&B_{1}\\ \end{array}\right],$
@@ 4 Main Results
The ultimate goal of this line of research is the computation of controllers whose factorized representation possesses a desired sparsity pattern or which inherits the structure of a closely related factorization, from which we may shift by employing numerically reliable procedures. This would allow us, for instance, to compute controllers that can be implemented as a “ring” network (recall Fig. 1) or as a “line” network, which is important for motion control of vehicles moving in “platoon” formation (see [25]). A key insight in this regard is the fact that the most prominent classical results in control theory, such as the Youla Parameterization, state the expression of a stabilizing controller via a coprime factorization over $\mathbb{S}$ of its TFM.
@@
A poignant insight into the relationship between the SRTR and NRF representations is hinted at in Fig. 2. As can be seen from the indicated diagram, an SRTR pair bears much stronger connections with a network’s state dynamics than its NRF counterpart. Indeed, it is this very affinity which serves to motivate the study of SRTR representations, in contrast to the more input-output oriented perspective offered by an NRF pair. This also makes SRTR pairs a far more appealing tool for system analysis than their correspondents in the NRF formalism, as the latter was primarily developed with the aim of controller synthesis (see [24] and [27]).
@@ 4.2 Getting from SRTR pairs to Left Coprime Factorizations
In this subsection, we show that any SRTR pair $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$ with both $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ having no poles located outside of $\mathbb{S}$ can produce an entire class of factorizations which are left coprime over $\mathbb{S}$. Notice that, for any SRTR pair $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$, the TFM $\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ is improper and has exactly $p$ poles at infinity of multiplicity one, and recall that the poles of $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ can be placed at will, as per Remark 3.14.
@@ 1. $\mathbf{a)}$
(20) $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{cc|cc}A_{x}-\lambda I_{p}&A_{12}&A_{x}-A_{11}+A_{12}K&B_{1}\\ O&A_{22}+KA_{12}-\lambda I_{n-p}&KA_{12}K+A_{22}K-KA_{11}-A_{21}&KB_{1}+B_{2}\\ \hline\cr U&O&U&O\\ \end{array}\right]\normalsize$
@@ 4.4 Stable and Stabilizing SRTR Implementations
In this subsection, we elaborate upon Remark 3.14 and present the means of implementing distributed controllers via their SRTR representations. The results given here are, in many ways, the analogues of Theorems III.4 and III.6 from [24], thus highlighting the connections being showcased in Fig. 2. Before doing so, we introduce the following definitions.
@@
Notice that the stabilizing and distributed controller from (22a) always has a stable TFM in the discrete-time case, as previously anticipated in Remark 3.14. Recalling the parallels discussed in Remark 3.9, this conclusion comes naturally and serves to further emphasize the close connections between the SRTR representations and the celebrated SLS framework [35], for discrete-time systems.
@@
In the continuous-time case, the sparsity structure of the distributed controller from (22a) is much more versatile than the ones obtained in [10]. We do not restrict $\mathbf{K}_{d}(\lambda)$ to possessing only banded nonzero entries and the generic sparsity structure it inherits from $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ offers more flexibility with respect to the network’s communication infrastructure. Note that, in contrast to the results from [10], we make no assumptions on the spatial invariance of the investigated networks.
@@
Let $\mathbf{G}\in\mathbb{R}(\lambda)^{m\times p}$ be given by a stabilizable and detectable realization and let $\mathbf{K}\in\mathbb{R}(\lambda)^{p\times m}$ be a stabilizing controller of $\mathbf{G}(\lambda)$. Let $\mathbf{K}(\lambda)$ be given by a minimal realization (4a)-(4b) and by a stable SRTR pair $(\mathbf{W}(\lambda),\mathbf{V}(\lambda))$ associated with the latter via (15). Let $e_{i}$ be the $i^{\text{th}}$ vector in the canonical basis of $\mathbb{R}^{p}$. By implementing stabilizable and detectable realizations for $e_{i}^{\top}\mathbf{K}_{d}(\lambda)$, with $i\in 1:p$, and by computing the command signals as in (22b), the state dynamics of the closed-loop system from Fig. 4 are made asymptotically stable (see Section 5.3 in [40]).
@@
For the means to obtain realizations of the type described in Theorem 4.21, consult Remark III.7 in [24]. Fortuitously, Remark III.8 in [24] also holds for block-row implementations (recall Remark 3.5) of the control laws from (22a)-(22b). Additionally, the synthesis techniques described in [24] and [27] can be employed in order to obtain the stabilizing controllers that are required by the results given in this subsection. Such a design scenario, in which the techniques from [24] and [27] are used to obtain a stabilizing controller which generates sparse SRTR-based control laws, will be showcased in Section 5 through a comprehensive numerical example.
@@
Among the chief benefits of the distributed control schemes proposed in Theorems 4.15 and 4.21 is that they provide stability guarantees in both the continuous- and discrete-time contexts. In contrast to this statement, we now refer to the proof of sufficiency for Theorem 2 from [35] and to the proof of Lemma 4 from [35]. The fact that $\frac{1}{z}I_{p}$ is a discrete-time stable TFM has been employed in these proofs so as to provide guarantees for both internal stability (recall Definition 4.13) and for the fact that the free evolution of the plant’s and the controller’s states tends to zero from any finite initial conditions (see the proof of Theorem 4.21 in Appendix D), in closed-loop configuration. Thus, the arguments made in [35] cannot be employed for the continuous-time setting, since $\frac{1}{s}I_{p}$ is not a stable TFM in the latter time domain.
@@
4.5 Forming structured and stabilizing STRT-based implementations
In light of Theorem 4.10 and of the results presented in Section III of [27], it is tempting to envision a procedure that produces stable SRTR pairs by way of structured LCFs over $\mathbb{S}$. However, a closer inspection of Theorem 4.21 suggests a different approach.
@@ 1. $iii)$
    $\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(KA_{11}-KA_{12}K+A_{21}-A_{22}K)e_{p,j}\left(1-e_{p,i}^{\top}\mathcal{B}_{\mathbf{W}}e_{p,j}\right)=O,\forall j\in 1:p$;
@@ 2. $iv)$
    $\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(KB_{1}+B_{2})e_{m,k}\left(1-e_{p,i}^{\top}\mathcal{B}_{\mathbf{V}}e_{m,k}\right)=O,\forall k\in 1:m$;
@@ 3. $v)$
    $\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}I_{n-p-n_{i}}&O\end{bmatrix}^{\top}=O$;
@@ 4. $vi)$
    $\left\{\lambda\in\mathbb{C}\ :\ \det\left(\lambda I_{n_{i}}-\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}^{\top}\right)=0\right\}\subseteq\mathbb{S}$.
@@ 1. $\mathbf{a)}$
(4.9a) $e_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=e_{p,i}^{\top}\begin{bmatrix}A_{11}-A_{12}K&B_{1}\end{bmatrix};$
@@
(4.9b) $e_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\begin{array}[]{c|c}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}^{\top}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\\ \hline\cr e_{p,i}^{\top}A_{12}Q_{i}^{\top}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}\phantom{}^{\top}&\widetilde{D}_{i}\\ \end{array}\right],$

### v7

@@ 3.4 System Response-Type Realizations
Let the system described by the realization (4a)-(4b), of order $n$, have a TFM be denoted $\mathbf{G}(\lambda)$. Then, a pair $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$ of proper TFMs, where $\mathbf{W}(\lambda)\in\mathbb{R}^{p\times p}(\lambda)$ and $\mathbf{V}(\lambda)\in\mathbb{R}^{p\times m}(\lambda)$, with $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ having a McMillan degree of at most npn-p and which satisfy
@@
It is worthwhile to point out that the two TFMs which make up the SRTR pair denoted $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$ fulfill the same role as the TFMs $\widetilde{\mathbf{R}}^{+}(z)$ and $\widetilde{\mathbf{N}}(z)$ from the state iteration equation of (18) from [SLS]. Denoting by $\mathbf{B}(z)$ the $\mathcal{Z}$-transform of the state vector belonging to the controller, the aforementioned state iteration is expressed in the same manner as (8) via the identity
@@
It is this very fact which brought about the designation of System Response-Type Realization and which can be exploited to obtain sparse and scalable representations for systems in the context of distributed control. Note that, in contrast to the implementations from [SLS], the one from (8) is also available for continuous-time systems.
@@
In the sequel, SRTR pairs of controllers will be employed to implement control laws in a distributed fashion. However, in contrast to the approach from [SLS] and its implementations of type (14), we point out that closed-loop maps are not intrinsically linked to the definition of an SRTR pair. Therefore, the newly introduced representation bears all the hallmarks of the so-called implementation matrices from [Separ], in which a mutation of the classical SLS paradigm allows for the decoupling of controller synthesis from closed-loop design. Although the framework proposed in the sequel shares this benefit, we continue to refer to SRTR pairs as such due to (14) being the inspiration for the proposed representation and, more importantly, due to not wishing to overshadow their significant potential for system analysis, which would inevitably result from naming them “implementation matrices”.
@@
Another feature of the SRTR representation is that any pair $\big(\hskip-0.7113pt\mathbf{W}\hskip-0.7113pt(\lambda),$ $\mathbf{V}(\lambda)\big)$ possesses the same sparsity pattern with its subsequent NRF pair $\big(\mathbf{\Phi}(\lambda),\mathbf{\Gamma}(\lambda)\big)$, obtained via (9a)-(9b). Therefore, $\mathbf{W}(\lambda)$ is lower triangular if and only if $\mathbf{\Phi}(\lambda)$ is lower triangular and, similarly, $\mathbf{V}(\lambda)$ is tridiagonal if and only if $\mathbf{\Gamma}(\lambda)$ is tridiagonal.
@@ 1. $\mathbf{a)}$
(15) $\small\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}=\left[\begin{array}[]{c|cc}A_{22}+KA_{12}-\lambda I_{n-p}&KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\\ \hline\cr A_{12}&A_{11}-A_{12}K&B_{1}\\ \end{array}\right],$
@@ 4 Main Results
The ultimate goal of this line of research is the computation of controllers whose factorized representation possesses a desired sparsity pattern or which inherits the structure of a closely related factorization, from which we may shift by employing numerically reliable procedures. This would allow us, for instance, to compute controllers that can be implemented as a “ring” network (recall Fig. 1) or as a “line” network, which is important for motion control of vehicles moving in “platoon” formation (see [plutonizare]). A key insight in this regard is the fact that the most prominent classical results in control theory, such as the Youla Parameterization, state the expression of a stabilizing controller via a coprime factorization over $\mathbb{S}$ of its TFM.
@@
A poignant insight into the relationship between the SRTR and NRF representations is hinted at in Fig. 2. As can be seen from the indicated diagram, an SRTR pair bears much stronger connections with a network’s state dynamics than its NRF counterpart. Indeed, it is this very affinity which serves to motivate the study of SRTR representations, in contrast to the more input-output oriented perspective offered by an NRF pair. This also makes SRTR pairs a far more appealing tool for system analysis than their correspondents in the NRF formalism, as the latter was primarily developed with the aim of controller synthesis (see [NRF] and [aug_sparse]).
@@ 4.2 Getting from SRTR pairs to Left Coprime Factorizations
In this subsection, we show that any SRTR pair $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$ with both $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ having no poles located outside of $\mathbb{S}$ can produce an entire class of factorizations which are left coprime over $\mathbb{S}$. Notice that, for any SRTR pair $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$, the TFM $\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}$ is improper and has exactly $p$ poles at infinity of multiplicity one, and recall that the poles of $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ can be placed at will, as per Remark 3.14.
@@ 1. $\mathbf{a)}$
(20) $\begin{bmatrix}\mathbf{M}(\lambda)&\hskip-5.69054pt\mathbf{N}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{cc|cc}\hskip-2.84526ptA_{x}-\lambda I_{p}&A_{12}&A_{x}-A_{11}+A_{12}K&B_{1}\\ O&\hskip-5.69054ptA_{22}+KA_{12}-\lambda I_{n-p}&KA_{12}K+A_{22}K-KA_{11}-A_{21}&\hskip-5.69054ptKB_{1}+B_{2}\\ \hline\cr U&O&U&O\\ \end{array}\right]\normalsize$
@@ 4.4 Stable and Stabilizing SRTR Implementations
In this subsection, we elaborate upon Remark 3.14 and present the means of implementing distributed controllers via their SRTR representations. The results given here are, in many ways, the analogues of Theorems III.4 and III.6 from [NRF], thus highlighting the connections being showcased in Fig. 2. Before doing so, we introduce the following definitions.
@@
Notice that the stabilizing and distributed controller from (22a) always has a stable TFM in the discrete-time case, as previously anticipated in Remark 3.14. Recalling the parallels discussed in Remark 3.9, this conclusion comes naturally and serves to further emphasize the close connections between the SRTR representations and the celebrated SLS framework [SLS], for discrete-time systems.
@@
In the continuous-time case, the sparsity structure of the distributed controller from (22a) is much more versatile than the ones obtained in [Explicit]. We do not restrict $\mathbf{K}_{d}(\lambda)$ to possessing only banded nonzero entries and the generic sparsity structure it inherits from $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ offers more flexibility with respect to the network’s communication infrastructure. Note that, in contrast to the results from [Explicit], we make no assumptions on the spatial invariance of the investigated networks.
@@
Let $\mathbf{G}\in\mathbb{R}(\lambda)^{m\times p}$ be given by a stabilizable and detectable realization and let $\mathbf{K}\in\mathbb{R}(\lambda)^{p\times m}$ be a stabilizing controller of $\mathbf{G}(\lambda)$. Let $\mathbf{K}(\lambda)$ be given by a minimal realization (4a)-(4b) and by a stable SRTR pair $(\mathbf{W}(\lambda),\mathbf{V}(\lambda))$ associated with the latter via (15). Let $e_{i}$ be the $i^{\text{th}}$ vector in the canonical basis of $\mathbb{R}^{p}$. By implementing stabilizable and detectable realizations for $e_{i}^{\top}\mathbf{K}_{d}(\lambda)$, with $i\in 1:p$, and by computing the command signals as in (22b), the state dynamics of the closed-loop system from Fig. 4 are made asymptotically stable (see Section 5.3 in [zhou]).
@@
For the means to obtain realizations of the type described in Theorem 4.21, consult Remark III.7 in [NRF]. Fortuitously, Remark III.8 in [NRF] also holds for block-row implementations (recall Remark 3.5) of the control laws from (22a)-(22b). Additionally, the synthesis techniques described in [NRF] and [aug_sparse] can be employed in order to obtain the stabilizing controllers that are required by the results given in this subsection. Such a design scenario, in which the techniques from [NRF] and [aug_sparse] are used to obtain a stabilizing controller which generates sparse SRTR-based control laws, will be showcased in Section 5 through a comprehensive numerical example.
@@
Among the chief benefits of the distributed control schemes proposed in Theorems 4.15 and 4.21 is that they provide stability guarantees in both the continuous- and discrete-time contexts. In contrast to this statement, we now refer to the proof of sufficiency for Theorem 2 from [SLS] and to the proof of Lemma 4 from [SLS]. The fact that $\frac{1}{z}I_{p}$ is a discrete-time stable TFM has been employed in these proofs so as to provide guarantees for both internal stability (recall Definition 4.13) and for the fact that the free evolution of the plant’s and the controller’s states tends to zero from any finite initial conditions (see the proof of Theorem 4.21 in Appendix D), in closed-loop configuration. Thus, the arguments made in [SLS] cannot be employed for the continuous-time setting, since $\frac{1}{s}I_{p}$ is not a stable TFM in the latter time domain.
@@
4.5 Forming structured and stabilizing SRTR-based implementations
In light of Theorem 4.10 and of the results presented in Section III of [aug_sparse], it is tempting to envision a procedure that produces stable SRTR pairs by way of structured LCFs over $\mathbb{S}$. However, a closer inspection of Theorem 4.21 suggests a different approach.
@@ 1. $iii)$
    $\begin{bmatrix}O&\hskip-5.69054ptI_{n_{i}}\end{bmatrix}\hskip-2.84526ptQ_{i}(KA_{11}-KA_{12}K+A_{21}-A_{22}K)e_{p,j}\left(1-e_{p,i}^{\top}\mathcal{B}_{\mathbf{W}}e_{p,j}\right)\hskip-2.84526pt=O,\forall j\in 1:p$;
@@ 2. $iv)$
    $\begin{bmatrix}O&\hskip-5.69054ptI_{n_{i}}\end{bmatrix}\hskip-2.84526ptQ_{i}(KB_{1}+B_{2})e_{m,k}\left(1-e_{p,i}^{\top}\mathcal{B}_{\mathbf{V}}e_{m,k}\right)=O,\forall k\in 1:m$;
@@ 3. $v)$
    $\begin{bmatrix}O&\hskip-5.69054ptI_{n_{i}}\end{bmatrix}\hskip-2.84526ptQ_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}I_{n-p-n_{i}}&O\end{bmatrix}^{\top}=O$;
@@ 4. $vi)$
    $\left\{\lambda\in\mathbb{C}\ :\ \det\left(\lambda I_{n_{i}}-\begin{bmatrix}O&\hskip-5.69054ptI_{n_{i}}\end{bmatrix}Q_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}O&\hskip-5.69054ptI_{n_{i}}\end{bmatrix}^{\top}\right)=0\right\}\subseteq\mathbb{S}$.
@@ 1. $\mathbf{a)}$
(4.9a) $\hskip-2.84526pte_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=e_{p,i}^{\top}\begin{bmatrix}A_{11}-A_{12}K&B_{1}\end{bmatrix}$;
@@
(4.9b) $e_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\begin{array}[]{c|c}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}Q_{i}(A_{22}+KA_{12})Q_{i}^{\top}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}^{\top}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\\ \hline\cr e_{p,i}^{\top}A_{12}Q_{i}^{\top}\begin{bmatrix}O&I_{n_{i}}\end{bmatrix}\hskip-2.84526pt\phantom{}^{\top}&\widetilde{D}_{i}\\ \end{array}\right],$